# Bridging LLMs and KGs without Fine-Tuning: Intermediate Probing Meets Subgraph-Aware Entity Descriptions

### v3

@@ Abstract.
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). By contrast, Large Language Models (LLMs) encapsulate extensive world knowledge and exhibit powerful context modeling capabilities, making them promising for mitigating the limitations of traditional methods. However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance. In this work, we propose a novel framework that synergizes the strength of LLMs with robust knowledge representation to enable effective and efficient KGC. We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances. These represenation are then utilized to train a data-effecient classifier tailored specifically for KGC tasks. To bridge the semanitic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions, ensuring that the representations are comprehensive and well aligned with the KG structure. Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach. We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by 188×188$\times$188 × and accelerating training and inference by 26.11×26.$1$.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019). By leveraging the powerful contextual modeling capabilities and extensive stored world knowledge of LLMs, we can effectively mitigate these challenges, and enhance the performance of KGC (Wei et al., 2024).
These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2024), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2024), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2024). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
Answering this question necessitates a deeper look at the limitations of LLMs compared to their fine-tuned counterparts. Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023). Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages, without explicitly fine-tuning on such tasks (Li et al., 2024).
In this paper, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1. Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC. To bridge the gap between LLMs and KGs, we incorporate an innovative subgraph sampling technique to generate model-friendly entity descriptions, enabling the LLMs to better interpret and disambiguate relational patterns to further enhance knowledge representation. By focusing on these intermediate LLM representations, our framework mitigates hallucination risks while retaining the benefits of a non-fine-tuned paradigm, allowing for both high accuracy and reduced computational costs.
Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs. Our approach enhances GPU memory efficiency by a factor of 188×188$\times$188 × and accelerates both training and inference processes by 26.11×26.$1$. Our contributions can be summarized as follows:
@@
  • We propose, to the best of our knowledge, the first framework that directly harnesses the intermediate representation space of non-fine-tuned LLMs for KGC tasks. To bridge the semantic gap between the LLM and the KG, we design a subgraph-based entity description generator, ensuring that the model’s internal understanding of entities aligns effectively with KG structures.
  • We provide an in-depth analysis of why fine-tuned LLMs typically achieve superior KGC performance, identifying the core factors responsible for these gains. Building on these insights, we present novel optimization strategies for non-fine-tuned LLMs and are the first to enable them to achieve KGC performance comparable to their fine-tuned counterparts.
@@
  • We perform extensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method. These substantial improvements highlight the scalability and practical applicability of our method, making it a viable solution for large-scale KGC applications where computational resources and time are critical constraints.
@@ 2. Method
Our method is proposed in face of the limitations of non-fine-tuned LLMs on KGC. In Section 2.2, we present how to precisely extract context-aware representation of knowledge from intermediate layers of LLMs. In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling.
@@ 2.1. Formulation and Notations
A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$. It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the textual description of each entity and each relation.
@@ 2.2. Extracting Knowledge Representations
Knowledge graphs are composed of numerous triples that encapsulate complex relational and semantic information. For non-fine-tuned LLMs, interpreting these intricate structures poses significant challenges, often leading to the retrieval of irrelevant information and the generation of incorrect or misleading responses.
To enhance the performance of LLMs in KGC, our method leverages the internal judgment mechanisms of LLMs, as shown in (Zou et al., 2023). Specifically, we strategically harness the intermediate layers of LLMs to extract rich, context-aware knowledge representations that are pivotal for accurate KGC.
@@
First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers. For triple classification, where the objective is to determine the plautibility of triples in the KG, we employ a prompt template, $PT_{1}$, that creates positive and negative sample pairs for training, as illustrated in Figure 2. Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
@@
Thereafter, we use existing triples in the knowledge graph to construct positive samples $S^{+}$, and employ the negative sampling technique (Lv et al., 2022) to obtain negative samples $S^{-}$. This results in a dataset obtaining $N$ training samples. Each training sample is input into the LLM to obtain representations from its intermediate layers.
Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
(1) $Rep_{j}^{i}\left(s^{+}\right)=\{M_{1:i}\left(s^{+}\right)\left[j\right],0\textless i\textless L,0\textless j\textless N\},$
@@
where $M$ represents the language model, $M_{1:i}$ represents the first $i$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$. Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models. Negative samples are processed similarly to ensure consistency in representation extraction.
@@
The extracted representations $Rep_{j}^{i}(s^{+})$ and $Rep_{j}^{i}(s^{-})$ serve as input features for training a lightweight classifier to distinguish between plausible and implausible triples by learning a decision boundary that effectively separates the hidden states of positive samples from those of negative samples. Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
@@
Figure 2. Prompt templates to stimulate the LLM.
@@ 2.3. Subgraph Entity Description Generator
The semantics of entities in a KG can be complex and ambiguous, posing significant challenges for LLMs to interpret accurately based solely on entity names. To overcome these challenges and enhance the semantic comprehension of entities by LLMs without resorting to computationally intensive fine-tuning, we introduce the subgraph entity description generator, which enriches entity representations by providing additional contextual descriptions derived from their local subgraph structures. Specifically, we extend the prompt template constructed in Section 2.2. We focus on aligning the language model’s internal entity understanding with those in the knowledge graph. Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2. By introducing the entity description $D(e)$, LLMs are able to generate more precise and distinguishable hidden states, thereby improving the overall classification performance.
Most KGs contain entity descriptions that can be directly used as $D(e)$. However, to address incomplete entity descriptions within certain KG datasets, we propose a method to generate $D(e)$. Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more precise $D(e)$ by incorporating subgraph information in two distinct ways. For the first approach, we concatenate all the triples within the one-hop subgraph around the entity $e$, thereby forming $D(e)$ (Zhang et al., 2024). However, providing information in the form of explicit rules may be less effective for LLM to retrieve and leverage knowledge (Bian et al., 2024), as the structure of these subgraph triples may significantly differ from the text seen during the pre-training phase. Therefore, our second approach leverages a language model to transform these triples into a narrative format that is more comprehensible to the LLM. By rephrasing the triples from the one-hop subgraph around the entity $e$ into natural language text, we produce a more refined and model-friendly version of $D(e)$ that is more effective for LLMs to retrieve and leverage knowledge. Specifically:
(2) $D(e)=\operatorname{Rephrase}(\{\operatorname{Verbalize}(\{h_{i}\}\{r_{i}\}\{t_{i}\}),$
$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}),$
@@
where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); For the first approach, $\operatorname{Rephrase(\cdot)}$ is a string concatenation function that joins each verbalized triple statement with a separator. For the second approach, we employ a LLM with the in-context learning technique, then ask it to generate model-friendly descriptions of entities based on the verbalized triples in their subgraph.
By leveraging the subgraph entity description generator, we align the LLM’s internal entity understanding with those in the KG, akin to entity linking process. This alignment not only mitigates ambiguities but also enhances LLM’s semantic comprehension, ultimately boosting KGC performance.

### v4

# Unlock the Power of Frozen LLMs in Knwoledge Graph Completion

@@ Abstract.
Traditional knowledge graph completion (KGC) methods rely solely on structural information, struggling with the inherent sparsity of knowledge graphs (KGs). By contrast, Large Language Models (LLMs) encapsulate extensive world knowledge and exhibit powerful context modeling capabilities, making them promising for mitigating the limitations of traditional methods. However, direct fine-tuning of LLMs for KGC, though effective, imposes substantial computational and memory overheads, while utilizing non-fine-tuned LLMs is efficient but yields suboptimal performance. In this work, we propose a novel framework that synergizes the strengths of LLMs with robust knowledge representation to enable effective and efficient KGC. We extract the context-aware hidden states of knowledge triples from the intermediate layers of LLMs, thereby capturing rich semantic and relational nuances. These representations are then utilized to train a data-efficient classifier tailored specifically for KGC tasks. To bridge the semantic gaps between LLMs and KGs, we employ subgraph sampling on KGs to generate model-friendly entity descriptions. We further adopt sliced mutual information (SMI) as a principled metric to quantify the task-specific information encoded in these representations. Extensive experiments on standard benchmarks validate the efficiency and effectiveness of our approach. We achieve a 47% relative improvement over previous methods based on non-fine-tuned LLMs and, to our knowledge, are the first to achieve classification performance comparable to fine-tuned LLMs while enhancing GPU memory efficiency by $188\times$ and accelerating training and inference by $26.11\times$.
Knowledge Graph Completion, Large Language Models
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: CIKM; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Reasoning about belief and knowledgeccs: Computing methodologies Information extractionccs: Computing methodologies Neural networksccs: Computing methodologies Supervised learning by classification
@@ 1. Introduction
Knowledge graph completion has become a crucial endeavor to improve the application of knowledge graphs (Chen et al., 2020), which aims to complete the knowledge graph and expand its scale by predicting the potential relationship between existing entities and discovering new relational facts (Shen et al., 2022b). Traditional knowledge graph embedding methods only utilize structure information from the knowledge graph (Bordes et al., 2013; Trouillon et al., 2016; Sun et al., 2019), which is vulnerable to the sparseness of KGs (Yao et al., 2019). Leveraging the powerful contextual modeling capabilities and the implicit world knowledge encoded in LLMs, recent works have demonstrated the potential to effectively mitigate these issues and enhance KGC performance (Wei et al., 2024).
These approaches generally fall into two types: using non-fine-tuned LLMs with prompting techniques and fine-tuned LLMs specifically for KGC. The first type leverages LLMs with prompt techniques, such as in-context learning. These techniques enable models to perform KGC directly through natural language outputs (Yang et al., 2024; Wei et al., 2024; Yao et al., 2025), harnessing the pre-trained knowledge embedded within these models without additional training. While this approach offers computational efficiency, it may not fully capture the task-specific nuances required for optimal KGC performance. In contrast, the latter type involves fine-tuning the LLMs using supervised techniques to adapt models specifically for KGC tasks. This method involves fine-tuning LLMs, such as LLaMA, through the use of training samples (Yao et al., 2025), which can lead to improved performance by aligning the models more closely with the KGC objectives (Zhang et al., 2024; Yao et al., 2025). However, fine-tuning demands significant increases in time and resource expenditure. This dichotomy between computational efficiency and task-specific performance raises a fundamental question: How can we harness LLMs for effective and efficient KGC without incurring the high costs associated with fine-tuning?
Answering this question necessitates a deeper understanding of why fine-tuned LLMs outperform their non-fine-tuned counterparts. Non-fine-tuned LLMs, while cost-effective, risk generating misleading outputs in the form of hallucinations (Zhao et al., 2023) and underutilizing their latent knowledge (Wang et al., 2023). Moreover, the intricate relational structures and semantic subtleties embedded in KGs compound the challenge, since LLMs are not inherently designed to parse or disambiguate complex entity linkages (Li et al., 2024). Crucially, fine-tuning mitigates hallucinations (Zhou et al., 2024) and aligns LLM representations with KG semantics through supervised parameter updates.
Motivated by this insight, we propose a novel framework that reconciles these competing demands by fully exploiting the knowledge representation capabilities of LLMs without fine-tuning, as shown in Figure 1. Our method strategically extracts intermediate hidden states from LLMs to capture rich, context-aware representations of knowledge triples, which are then used to train a lightweight, data-efficient classifier specifically tailored for KGC. To better align LLMs with KG semantics, we introduce a novel subgraph sampling technique that generates model-friendly entity descriptions, facilitating the interpretation and disambiguation of relational patterns and thereby enhancing knowledge representation. Furthermore, we introduce Sliced Mutual Information (SMI) as a principled information-theoretic measure to quantify the task-specific information encoded within these representations. By systematically integrating these techniques, our framework effectively reduces hallucinations and semantic misalignment commonly observed in frozen LLMs, achieving accuracy comparable to fine-tuned counterparts while significantly reducing computational overhead.
Extensive experiments demonstrate that our framework not only outperforms existing non-fine-tuned LLM-based methods by $47\%$ relative improvement but also achieves classification performance comparable to fine-tuned LLMs. Our approach enhances GPU memory efficiency by a factor of $188\times$ and accelerates both training and inference processes by $26.11\times$. Our contributions can be summarized as follows:
@@
  • We propose, to the best of our knowledge, the first framework that directly leverages the intermediate representation space of frozen LLMs for KGC tasks. To facilitate semantic alignment between LLM representations and KG entities, we introduce a novel subgraph-based entity description generator, providing LLMs with rich, model-friendly textual contexts to effectively bridge the semantic gap.
  • We provide an in-depth analysis to reveal the underlying reasons behind the superior performance of fine-tuned LLMs. We further employ SMI to quantify task-specific information encoded in LLM representations. Guided by these insights, we propose novel optimization strategies that, for the first time, enable frozen LLMs to match the KGC performance of their fine-tuned counterparts.
@@
  • We conduct comprehensive experiments across standard KGC datasets to validate the effectiveness and efficiency of our proposed method. These substantial improvements highlight the scalability and practicality of our method, making it a viable solution for large-scale KGC applications under resource-constrained scenarios.
@@ 2. Method
Our method is proposed in face of the limitations of non-fine-tuned LLMs on KGC. In Section 2.2, we present how to precisely extract context-aware representation of knowledge from intermediate layers of LLMs. In Section 2.3 we tackle complex relationships in KGs, by the generation of model-friendly entity descriptions through subgraph sampling. Furthermore, in Section 2.4, we employ SMI as a principled information-theoretic metric to quantify the task-specific information encoded within the extracted representations, providing deeper insights into their effectiveness.
@@ 2.1. Formulation and Notations
A knowledge graph, $G$, comprises a set of entities, $E$, and relations, $R$. It can be represented as a set of triples $G=\{(h,r,t)\}$, where $h\in E$ is the head entity, $t\in E$ is the tail entity, we collectively refer to the head entity and tail entity as $e$, and $r\in R$ represents the relation between them. $D$ is the set of long descriptions for each entity and relation. We denote $D(e)$, $D(r)$ as the long textual description of each entity and each relation (hereafter simply referred to as descriptions). Additionally, we define $EN(e)$ as the entity name of an entity $e$. For instance, given the entity ‘00952182’, its entity name is $EN(\text{`00952182'})=\text{`voice'}$, and its description $D(\text{`00952182'})$ can be ‘utter with vibrating vocal chords’.
@@ 2.2. Extracting Knowledge Representations
Figure 2. Prompt templates to stimulate the LLM.
Knowledge graphs comprise numerous triples encoding complex relational and semantic information. However, directly prompting pre-trained, frozen LLMs, despite their extensive world knowledge, frequently results in hallucinations or incorrect responses. Although fine-tuning effectively mitigates this issue by optimizing the model’s parameters to produce KG-consistent outputs, it incurs significant computational overhead. Motivated by recent findings (Zou et al., 2023; Orgad et al., 2024) that accurate relational knowledge is often internally encoded within LLMs yet not reliably reflected in their outputs, we propose directly probing intermediate layer representations as a principled alternative to fine-tuning. This approach leverages the internal semantic encoding capabilities of frozen LLMs, thereby mitigating the computational overhead and hallucination problems associated with direct textual generation.
@@
First, we construct specialized prompt templates designed to stimulate the LLM to generate the necessary classification information within its intermediate layers. For triple classification, where the objective is to determine the plausibility of KG triples, we employ a prompt template $PT_{1}$ to construct positive and negative sample pairs for training, as illustrated in Figure 2. Notably, the prompt avoids explicitly describing the positivity or negativity of triples.
@@
Subsequently, we construct positive samples $S^{+}$ from existing KG triples, and employ the negative sampling technique (Yao et al., 2019) to generate negative samples $S^{-}$. This process results in a training dataset containing $N$ samples. Each training sample is tokenized and fed into the LLM to obtain representations from its intermediate layers. Specifically, we capture plausibility information and extract hidden states for positive samples $s^{+}\in S^{+}$ as follows:
(1) $Rep_{j}^{l}\left(s^{+}\right)=\{M_{1:l}\left(s^{+}\right)\left[j\right],0\textless l\textless L,0\textless j\textless N\},$
@@
where $M$ represents the language model, $M_{1:l}$ represents the first $l$ layers of the model, $[j]$ refers to the hidden state corresponding to the $j$th token, $L$ and $N$ represent the number of layers in the model and the number of tokens resulting from tokenizing $s^{+}$, respectively. Consistent with findings in (Zou et al., 2023), we utilize the representation of the last token in each input text, as it is more conducive to information extraction due to the autoregressive nature of causal language models. Negative samples are processed similarly to ensure consistency in representation extraction.
@@
The extracted representations $Rep_{j}^{l}(s^{+})$ and $Rep_{j}^{l}(s^{-})$ serve as input features to train a lightweight classifier. This classifier distinguishes plausible from implausible triples by learning a decision boundary that effectively separates positive and negative samples in the representation space. Although the classifiers employed—such as multilayer perceptrons (MLPs) and logistic regression models—are inherently simple, their effectiveness is significantly amplified by the quality and richness of the intermediate layer representations, thereby enhancing the ability to make accurate predictions without the computational overhead associated with fine-tuning the entire LLM.
@@
Figure 3. Prompt used for generating model-friendly entity descriptions via LLM rephrasing. In-context examples omitted for brevity.
@@ 2.3. Subgraph Entity Description Generator
Entities in a KG often exhibit semantic ambiguity, posing significant challenges for frozen LLMs that rely solely on entity names. Fine-tuning alleviates this ambiguity by aligning LLM entity representations with KG entities. However, for frozen LLMs, achieving such semantic alignment without parameter updates remains challenging. To bridge this gap, we propose a principled subgraph-besed approach that generates rich, context-aware entity descriptions from local subgraph contexts. These descriptions provide rich semantic context, enabling frozen LLMs to achieve semantic alignment comparable to fine-tuned models, without costly parameter updates. Specifically, we achieve this by extending the prompt template constructed in Section 2.2. We focus on aligning the language model’s internal entity understanding with those in the KG. Therefore, we construct new prompts $PT_{2}$ and $PT_{4}$, as shown in Figure 2.
Most KGs do not contain explicit entity descriptions that can be directly used as $D(e)$, or provide only limited and incomplete descriptions. Therefore, we propose a method to generate comprehensive and informative entity descriptions $D(e)$. Unlike the approach in (Yang et al., 2024), which generates the entity description using only a single triple, our method constructs a more comprehensive and precise $D(e)$ by incorporating richer subgraph contexts in two distinct ways. For the first approach, following (Zhang et al., 2024), we construct $D_{1}(e)$ by concatenating the verbalized triples within the one-hop subgraph around the entity $e$:
(2) $D_{1}(e)=\operatorname{CONCAT}\big{(}\{\operatorname{Verbalize}(EN(h_{i}),EN(r_{i}),EN(t_{i})),$
$(h_{i},r_{i},t_{i})\in\operatorname{Subgraph}(e)\}\big{)},$
@@
where $\operatorname{Subgraph(\cdot)}$ is a function that automatically retrieves a specified number of triples in the one-hop subgraph where the entity resides; $\operatorname{Verbalize(\cdot)}$ is a function that transform the triple into its textual string, in this work, we use the linear verbalization (Baek et al., 2023); and $\operatorname{CONCAT(\cdot)}$ joins each verbalized triple statement with a separator.
However, presenting information explicitly as structured triples may hinder effective knowledge retrieval by LLMs (Bian et al., 2024), since the structure of these subgraph triples substantially deviates from the textual data encountered during pre-training. Motivated by this limitation, we propose a second, more model-friendly approach, denoted as $D_{2}(e)$, which employs a language model to rephrase these verbalized triples into a narrative format that is more comprehensible to the LLM. Specifically, we employ an in-context learning approach, where a small set of carefully selected examples are prepended to the prompt to guide the LLM to generate high-quality, narrative-like descriptions. For brevity, Figure 3 illustrates only the core instruction of our prompt, omitting the in-context examples. This approach produces more refined and model-friendly entity descriptions, enabling the LLM to more effectively retrieve and leverage knowledge during the KGC task.
By leveraging the subgraph entity description generator, we align the LLM’s internal understanding of entities with their counterparts in the KG, similar to an entity linking process. This approach addresses semantic ambiguity, achieving alignment comparable to fine-tuning while significantly reducing computational overhead.
2.4. Representation Evaluation via Sliced Mutual Information
To quantitatively assess the task-specific information encoded in the frozen-LLM representations, we adopt sliced mutual information (SMI), which scales more effectively to high-dimensional data compared to mutual information (Goldfeld and Greenewald, 2021; Wongso et al., 2023). Let the random variable $X$ denote the representations extracted from the intermediate layers, where $x_{k}=Rep_{j}^{l}(s_{k})\in\mathbb{R}^{d}$ corresponds to the representation of the $k$-th sample $s_{k}$. Let the random variable $Y$ denote the corresponding discrete KGC labels. Since labels $Y$ are discrete, the SMI is defined as the expectation of mutual information between the labels $Y$ and one-dimensional random projections of $X$, where these projections are sampled uniformly at random from the $d$-dimensional unit sphere (denoted $\mathbb{S}^{d-1}$). We estimate SMI via Monte-Carlo sampling combined with a scalar MI estimator (Kraskov et al., 2004):
$\widehat{SMI}(X;Y)\;=\;\frac{1}{m}\sum_{i=1}^{m}\;\hat{I}\big{(}\{\theta_{i}^{\top}x_{k}\}_{k=1}^{n};\{y_{k}\}_{k=1}^{n}\big{)},\quad\theta_{i}\sim\text{Unif}(\mathbb{S}^{d-1}).$
[…]
3. Experiments
3.1. Datasets
Dataset $|\mathbf{E}|$ $|\mathbf{R}|$ # Train # Valid # Test
FB13 75,043 13 316,232 5,908 23,733
WN11 38,696 11 112,581 2,609 10,544
FB15K-237N 13,104 93 87,282 14,082 16,452
WN18RR 40,943 11 86,835 6,068 6,268
UMLS 135 46 5,216 1,304 1,322
YAGO3-10 123,182 37 1,079,040 5,000 5,000
Table 1. Statistical information of datasets.
In our experiments, we employ five widely-used KG datasets for triple classification: FB13 (Socher et al., 2013), WN11 (Socher et al., 2013), FB15K-237N (Lv et al., 2022), WN18RR (Dettmers et al., 2018), and UMLS (Yao et al., 2019). For relation prediction, we adopt the YAGO3-10 dataset (Dettmers et al., 2018), and for entity prediction, we use WN18RR. Specifically, WN11 and WN18RR are subsets of WordNet, while FB13 and FB15K-237 (Toutanova and Chen, 2015) are subsets of Freebase. The FB15K-237N dataset is a modified version of FB15K-237. Additionally, UMLS is a medical semantic network comprising semantic types (entities) and semantic relations.
To address incomplete or low-quality entity descriptions in some KGs, we propose two systematic approaches to generate entity descriptions from local subgraph structures, as detailed in Section 2.3 (hereafter referred to as generated entity descriptions). Specifically, for the first approach in Section 2.3, we construct entity descriptions by concatenating ten randomly sampled triples from each entity’s one-hop subgraph, employing linear verbalization (Baek et al., 2023); we denote these as generated descriptions (Tri). For the second approach, we further utilize GPT-3.5-turbo to rephrase these concatenated triples into coherent, model-friendly descriptions via in-context learning; we denote these as generated descriptions (GPT).
To assess the quality and effectiveness of our generated entity descriptions, we also utilize existing entity descriptions (provided by the original datasets or external authoritative sources) as reference baselines, hereafter referred to as non-generated descriptions. Specifically, for FB13, entity descriptions are obtained from corresponding Wikipedia; for FB15K-237N, we adopt the descriptions provided by (Xie et al., 2016); for WN18RR, we utilize synset definitions from (Yao et al., 2019); and for UMLS, we employ the entity descriptions included in the original dataset.